OpenStudy (anonymous):

Given the function f(x) = 2(x + 8), find x if f(x) = 12

OpenStudy (anonymous):

You are given with f(x) = 12 So just plug in the value of f(x) \[2( x + 8) = 12\] Now divide by 2 both the sides and tell me what did you get...

OpenStudy (anonymous):

i got 40

OpenStudy (anonymous):

wait... :P never mind i did it wrong

OpenStudy (anonymous):

no.. It is not the values of x it is value of f(x) be careful..

OpenStudy (anonymous):

See the difference: f(x) = 12 Here the value of function itself is 12 f(12) Here the value of x is 12..

OpenStudy (anonymous):

Are you solving it or not @vetdoc

OpenStudy (anonymous):

would i subtract the12 by 8?

OpenStudy (anonymous):

Firstly divide by 2 both the sides..

OpenStudy (anonymous):

Show your steps how are you doing it and then I can tell what step is to make...

OpenStudy (anonymous):

2(x+8)=12 -- -- 2 2 x+8=6 -8 -8 x=-2

OpenStudy (anonymous):

Yes you are absolutely right.. Cheers...!